Find the effective yield, to the nearest hundreth percent, of an investment at 6% coumpounded continuesly?
Answer by stanbon(48535)

Find the effective yield, to the nearest hundreth percent, of an investment at 6% coumpounded continuesly?
---------------------
A(t) = Aoe^(0.06t)
A(t) = Ao*(e^0.06)^t
A(t) = Ao*1.061837^t
A(t) = Ao(1 + 0.06183)^t
--------------------------
The effective interest rate is 6.183%
================================
Cheers,
Stan H.